OpenStudy (anonymous):

Can some one please help me? I've already tryed to figure it out on my own..but its no use! A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air? Could you show me step by step so I understand it!? Thank you <3

OpenStudy (anonymous):

[(x + 40) mph](5 hours) = [(x - 40) mph](6 hours) 5x + 200 = 6x - 240 x = 440 mph

OpenStudy (anonymous):

The two sides can be set equal to each other because the distance is the same and the units are "miles".

OpenStudy (anonymous):

"x + 40" is the combined speed on this first leg of the trip, and "x - 40" is the combined speed on the trip back.

OpenStudy (anonymous):

The equation is in "miles" because "hours" times "mph" gives "miles".

OpenStudy (anonymous):

ahhh. I think I kind of get it.

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @magenxoxo Let's not stop at a fuzzy understanding. Please ask some questions.

OpenStudy (anonymous):

I will try doing the next word problem on my own and run it by u to see if its correct?! :)

OpenStudy (anonymous):

The first key to understanding this is that if the plane is pushing itself the same each way, that is "x" mph and with a 40 mph tailwind, it goes x + 40 mph and into the wind it goes x - 40 mph

OpenStudy (anonymous):

The second key to understanding this is that rate times time = distance and that is the same both ways. So we get an equation in distance.

OpenStudy (anonymous):

yes.. I understand that part :) so we use the distance formula.. d=rt

OpenStudy (anonymous):

So, the distances are the same, and that is expressed by the rate of "x + 40" times the time of 5 hours, going there. Coming back we have the rate of "x - 40" times the time of 6 hours. And then we just solve for "x".

OpenStudy (anonymous):

As a by-product, though the problem doesn't ask for it. we also know the distance, and that is 2400 miles each way because: (440 + 40)(5) = 2400

OpenStudy (anonymous):

Just as (440 - 40)(6) = 2400 on the way back.

OpenStudy (anonymous):

So, the plane, relative to the ground (not the relative to the air), goes 480 mph going there and 400 mph coming back.

OpenStudy (anonymous):

So, that's about all I can add to the understanding of that problem unless you want to ask questions. Or we can go on to another problem.

OpenStudy (anonymous):

ok! thank youuu :) i'm actually working on another problem now.. I wanted to try and see if i could do it. but if i need help i'll post it here :)

OpenStudy (anonymous):

Two bicycles depart from Miami Beach going in opposite directions. The first bicycle is traveling at 10 miles per hour. The second bicycle travels at 5 miles per hour. How long does it take until the bikes are 45 miles apart?

OpenStudy (anonymous):

@blondie16

OpenStudy (anonymous):

sorry i dont remember how to do this lol! hmmm let me see @abb0t @hartnn @Hero

OpenStudy (anonymous):

haha. its all goooood :)

OpenStudy (anonymous):

wouldn't it be 50? bc 10x5=50.... ur talkin about the seccond question i posted? correct?

OpenStudy (anonymous):

wouldn't the equation be (10mph)t + (5mph)t???

OpenStudy (anonymous):

OOOOOH

OpenStudy (anonymous):

idk.. I'm confused

OpenStudy (anonymous):

uhm.. 3 hours? haha.... that seems wrongg

OpenStudy (anonymous):

2 hours and 25 min?

OpenStudy (anonymous):

ahh. yay :) I was right!

OpenStudy (anonymous):

Thank youuu!!!

OpenStudy (anonymous):

aha.. yeah.. ur right.